OpenStudy (anonymous):

how to obtain equation for cosh(3x) in terms of cosh(x) and sinh(x)

OpenStudy (accessdenied):

Have you learned of the sum identites for cosh and sinh? cosh(x+y) = cosh x cosh y + sinh x sinh y sinh(x+y) = sinh x cosh y + sinh y cosh x If you know those, you could probably just apply this twice; once for 3x = 2x + x, and then for 2x = x + x.